a car is moving at as speed of 15m/s.what is the time taken by it to cover the distance of 1.2km ?
1
Expert's answer
2014-07-21T11:47:53-0400
As we know,we need to use the formula S=V*T, where S – distance, V – speed and T – time.

That meansthat T = S/V. To calculate this we need to use the same dimensions. Therefore, we need to convert kilometers into meters:
1.2 km =1.2 * 1000 m = 1200 m. (because, as we know, 1km=1000m)
Now we canuse our formula: T=S/V = 1200/15 = 80 s.
We cantransform seconds into minutes (1 min = 60 s). And the answer: 1 min 20 s.